OpenStudy (anonymous):

it takes a runner 3h, 58min, 32s to run a marathon. if the distance of the marathon is 42.4km/hr, what is the average speed of the runner

OpenStudy (shane_b):

\[Speed=\frac{distance}{time}\]Your distance is 42.4km and the time is 3:58:32. To use the formula above simply convert the time to hours:\[3h + (\frac{58}{60})h+\frac{32s}{3600s/h}=\large \text{~}3.98h\]So you end up with:\[Speed = \frac{42.4km}{3.98h}=?? km/h\]